Cotton White, Inc., makes specialty clothing for chefs. The company reported the following costs for 2015: Factory rent Company advertising Wages paid to seamstresses Depreciation on salespersons vehicles Thread Utilities for factory Cutting room supervisors salary Presidents salary Premium quality cotton material Buttons Factory insurance Depreciation on sewing machines Wages paid to cutters S 37,000 25,600 75,500 30,800 1,030 23,400 30,700 76,100 40,300 750 19,800 6,400 51,900 Required: 1. Compute the cost of direct materials for Cotton White. 2. Compute the cost of direct labor for Cotton White 3. Compute the cost of manufacturing overhead for Cottion White. F1 F2 F3 F4 FS F6
ⓘU ezto.mheducation.com/hm.tpx 3. Compute the cost of manufacturing overhead for Cotton White 4. Compute the total manufacturing cost for Cotton White 5. Compute the prime cost for Cotton White 6. Compute the conversion cost for Cotton White. 7. Compute the total period cost for Cotton White 求: F1 F2 F3 F4 FS F6
Answer #1

SOLUTION:

1) Direct material cost: $40,300

Working: Premium quality cotton material: $40,300

2) Direct labor cost: $127,400

Working: Wages paid to seamstresses + Wages paid to cutters = 75,500 + 51,900 = 127,400

3) Manufacturing overhead: $119,080

Working:

Factory rent + Thread + Utilities for factory + Cutting room supervisor's salary + Buttons + Factory insurance + Depreciation on sewing machines

37,000 + 1,030 + 23,400 + 30,700 + 750 + 19,800 + 6,400

= 119,080

4) Total manufacturing cost: $286,780

Working: 40,300 + 127,400 + 119,080 = 286,780

5) Prime Cost

Working: 40,300 + 127,400 = 167,700

6) Conversion Cost: 246,480

Working: 127,400 + 119,080 = 246,480

7) Period cost: $132,500

Working: Company advertising + Depreciation on salespersons' vehicles + President’s salary

25,600 + 30,800 + 76,100

= $132,500